# Adapt-$\infty$: Scalable Lifelong Multimodal Instruction Tuning via Dynamic Data Selection

### v1

@@ Abstract
Visual instruction datasets from various distributors are released at different times and often contain a significant number of semantically redundant text-image pairs, depending on their task compositions (i.e., skills) or reference sources. This redundancy greatly limits the efficient deployment of lifelong adaptable multimodal large language models, hindering their ability to refine existing skills and acquire new competencies over time. To address this, we reframe the problem of Lifelong Instruction Tuning (LiIT) via data selection, where the model automatically selects beneficial samples to learn from earlier and new datasets based on the current state of acquired knowledge in the model. Based on empirical analyses that show that selecting the best data subset using a static importance measure is often ineffective for multi-task datasets with evolving distributions, we propose Adapt-$\infty$, a new multi-way and adaptive data selection approach that dynamically balances sample efficiency and effectiveness during LiIT. We first construct pseudo-skill clusters by grouping gradient-based sample vectors. Next, we select the best-performing data selector for each skill cluster from a pool of selector experts, including our newly proposed scoring function, Image Grounding score. This data selector samples a subset of the most important samples from each skill cluster for training. To prevent the continuous increase in the size of the dataset pool during LiIT, which would result in excessive computation, we further introduce a cluster-wise permanent data pruning strategy to remove the most semantically redundant samples from each cluster, keeping computational requirements manageable. We validate the effectiveness and efficiency of Adapt-$\infty$ over a sequence of various multimodal instruction tuning datasets with various tasks, including (Knowledge) VQA, multilingual, grounding, reasoning, language-only, and multi-image comprehension tasks. Training with samples selected by Adapt-$\infty$ alleviates catastrophic forgetting, especially for rare tasks, and promotes forward transfer across the continuum using only a fraction of the original datasets. [^1] Code is released at https://github.com/adymaharana/adapt-inf.
@@ 1 Introduction
Multimodal instruction tuning (Liu et al., 2023b; Zhang et al., 2023a; Liu et al., 2023a; Gan et al., 2024; Yoon et al., 2024) has been actively explored to enhance visual reasoning or the generation ability of Multimodal Large Language Models (MLLMs) (Zhu et al., 2023; Li et al., 2023a; Tang et al., 2023; Team et al., 2023; Munasinghe et al., 2023; Yu et al., 2024; Zhang et al., 2024) following user intents by training models on human or machine-generated multi-task visual instruction tuning datasets (Li et al., 2023c; Yin et al., 2023; Xu et al., 2024; Chen et al., 2024). While many distributors continue to release new high-quality instruction-tuning tasks and datasets, continually adapting large models to these massive datasets over time is prohibitively costly and inefficient. Given a pre-trained MLLM and the continuous expansion of the dataset pool with a stream of instruction-tuning datasets, as commonly observed in the research community today, the challenge lies in developing an ever-evolving, instruction-following MLLM in the most data- and computation-efficient manner. This research question poses a realistic, sustainable instruction tuning scenario for MLLMs, distinct from conventional continual learning (Zenke et al., 2017; Yoon et al., 2018; Van de Ven & Tolias, 2019), which focuses on learning a sequence of disjoint tasks. Specifically, at each time step, we assume a new multimodal instruction-tuning dataset is added to the existing training pool, which already contains previous datasets. The model aims to train on samples from this continually expanding dataset pool, a scenario we refer to as Lifelong Instruction Tuning (LiIT). This scenario is critical as recent datasets tend to be exhaustively large to cover a variety of skills in uni- and multimodal tasks.
@@
To build an efficient, lifelong-evolving MLLM, we introduce Adapt-$\infty$: Adaptive Multi-way Pruning, an efficient and dynamic multimodal data selection strategy. At each time step, Adapt-$\infty$ selects the most beneficial samples for the current model from the data pool, adapting to the model’s evolving knowledge and changing dataset distributions. This is crucial for LiIT, as sample importance shifts over time. Adapt-$\infty$ operates in two major steps: (1) Task-based clustering. When introducing a new dataset, we integrate it into the training data pool and create pseudo-task clusters using gradient vectors to represent data samples. (2) Cluster-wise data selection. We select influential samples from each cluster for model training. Motivated by the observation that different score functions (Paul et al., 2021; Toneva et al., ; Coleman et al., ) define sample importance differently, we propose a new multi-way data selection approach that chooses the best scoring function from a pool of experts based on its discriminativeness (measured by entropy). This approach is able to select a skill-balanced subset of highly influential samples from the current training data pool. As the size of the data pool continues to grow with time, the inference step for computing scores and data representations in Adapt-$\infty$, as in most existing data selection strategies, can become prohibitively expensive. Hence, to maintain a sufficiently diverse yet computationally manageable data pool, Adapt-$\infty$ performs an additional step: (3) Permanent data pruning that removes semantically redundant samples from the data pool at the end of each time step, thereby continually controlling its size during LiIT.
To better assess the influence of multimodal samples, we also propose a new scoring function, called the image grounding score (IG), that measures the relative change in sample perplexity when the model is grounded by visual information. This metric prioritizes samples that effectively utilize and improve the multimodal skills of the MLLM and serves as an effective data selector in Adapt-$\infty$.
@@
We design an experimental setup for this previously unexplored scenario of lifelong multimodal instruction tuning using the pre-trained LLaVA 1.5 (Liu et al., 2023a) model on a stream of five visual instruction tuning datasets. As discussed earlier, experience replay alone is insufficient to mitigate forgetting, however, selecting a random subset from the data pool of past and new datasets reduces the forgetting rate from 26% to nearly 2%. Score-based data-selection strategies largely fail in this setting due to their inability to select task-balanced data subsets from multi-task datasets. In contrast, Adapt-$\infty$ minimizes the forgetting to just 0.9% using only a fraction of the training data pool while promoting forward transfer of skills and consistently achieving $>$100% relative gains.
@@
Adapt-$\infty$ provides an intuitive framework for dynamic data selection in the temporal scenario. We conduct extensive ablations of Adapt-$\infty$ to find its best-performing settings. One of our significant findings is that hidden layer outputs represent the semantic component whereas gradient vectors represent the skill component of samples. Hence, gradient vectors are more effective at pooling samples into pseudo-task clusters (see examples in Figures 2 and 4). Further, we find that gradients from the middle layer of the model lead to best overall performance, suggesting that skill retention could be localized to a few layers in LLMs. We also find that zero-order gradients (Hinton, 2022; Sung et al., 2024) are promising and computationally cheaper alternatives to backpropagated gradients for pseudo-task clustering (See Table 5). Analysis of the skill-wise breakdown of performance reveals that language-only skills are the easiest to retain and improve, whereas multilingual multimodal skills exhibit significant forgetting in LiIT.
@@ 2 Related Work
Data selection has been explored in the form of coreset selection in many works (Welling, 2009; Chen et al., 2010; Feldman et al., 2011). Uncertainty/loss/error-based methods estimate the difficulty of a sample from model confidence (Swayamdipta et al., 2020) or its training dynamics (Toneva et al., ; Paul et al., 2021; Bachem et al., 2015). Zheng et al. (2023) address catastrophic accuracy drop at high pruning rates; Maharana et al. (2024) represent datasetd as undirected graphs and employ message passing to select the best subset. Gadre et al. (2024) investigate data selection for CLIP Radford et al. (2021) models. Evans et al. (2024) use learnability score (Mindermann et al., 2022) to accelerate training of CLIP models.
Figure 2: A: Sample distributions for different visual language tasks in M3IT (Li et al., 2023b) based on two importance scores, EL2N and entropy. B: t-SNE visualization of sample vectors based on their gradients and features. C: Histogram of Perplexity and Image Grounding scores. We visualize a few samples from M3IT with prompts (black) and ground-truth answers (blue).
@@ 4.4 Combined Permanent-Temporal Data Pruning by Removing Redundancy
The steps of Adapt-$\infty$ outlined in the previous sections enable the model to effectively select the most beneficial samples for each pseudo-task (i.e., skill) and train on them adaptively, based on the current model and evolving dataset distributions. However, as the data pool grows with the release of new instruction-tuning datasets, the computational burden of updating gradient-based sample vectors and ranking their importance increases. To keep the computational requirements manageable over time, we further implement a permanent data pruning strategy to iteratively reduce the size of the entire dataset pool. At the end of training at each timestep, we measure pairwise cosine similarities between samples within each cluster and prune those with maximum similarities, as they are highly redundant and contain overlapping knowledge. The similarity is computed in the semantic space which is well represented using hidden layer outputs of the model as shown in Figure 4 (Abbas et al., 2023; Sorscher et al., 2022). Larger clusters are prioritized for pruning to fit a predefined data pool budget $D$ until all clusters retain a uniform number of samples. We refer to the version of Adapt-$\infty$ with this combined permanent-temporal pruning as Lite-Adapt-$\infty$.
@@ 5.1 Experimental Setup
We conduct our experiments using the LLaVA 1.5 multimodal large language model (Liu et al., 2023a). It is trained on top of the Vicuna LLM (Chiang et al., 2023) using a corpus of approximately 600K image-text caption pairs for pretaining of the vision projectors from the pre-trained CLIP visual encoder ViT-L/14 (Radford et al., 2021). Further, it is trained on the LLaVA instruction tuning dataset consisting of 665K text-only and vision-language instances. We adopt LoRA finetuning (Hu et al., 2021) of the LLaVA-1.5-7B model with the recommended hyperparameters. [^2] https://github.com/haotian-liu/LLaVA/tree/main
@@
Adapt-$\infty$ Setup. The optimal value of $k$ in the pseudo-task clustering step is computed from a grid search over values of $k$ between 5 and 50, and selected absed on the WSS value of clusters. [^3] Within the sum of squares (WSS) is sum of squared distance between samples and their cluster centroids. In the score-based sample selection step, we use a bin size of 50 and discard the top and bottom 5% of samples for computing entropy as well as for CCS sampling, to remove outliers, low-quality samples, and uninformative data (Zheng et al., 2023). We use perplexity, image grounding (Section 3.2), EL2N (Paul et al., 2021) and entropy (Coleman et al., ) score functions for $\mathcal{S}$ throughout the paper.
@@ 5.2 Main Results
Sequential training leads to catastrophic forgetting of skills. Multiple skills learned at $t$=0 (LLaVA) are forgotten at $t$=1,2 upon training the model on datasets containing a different set of skills. The M3IT dataset ($t$=1) does not contain grounding tasks and results in large drops in performance for the same. Similarly, the MiniGPT4 dataset ($t$=2) predominantly contains high-quality captions and causes forgetting of all other skills. The MANTIS dataset ($t$=3) improves performance on the MMMU dataset because it contains training instances with documents, charts, and visualization images. At $t$=4, the LAMM dataset contains object detection and keypoint detection tasks that promote recovery of the referring comprehension skill learned at $t=0$, presumably due to a similarity in their non-textual output modalities. Finally, VisionFLAN ($t$=5) recovers performance on all skills except grounding and MMMU. Surprisingly, VisionFLAN also induces recovery of multilingual and unimodal skills (MMLU) despite not containing those tasks in its dataset composition. This method incurs nearly 26% forgetting across all timesteps and retains only 68% of its all-time high performances at the final timestep (see Table 2).
@@
The use of gradient vectors in Adapt-$\infty$ to pool samples into pseudo-task clusters before pruning ensures that all skills are well-represented at each time step resulting in the lowest forgetting rate i.e., 0.9%. Further, the multi-way score-based selection of important samples within those clusters promotes forward transfer of skills and results in $>$100% relative gain unlike any other method in Table 2. The relative gain increases by nearly 5% on doubling the data budget from 25K to 50K samples and shows signs of plateauing with further increase in data.
@@ 6 Analysis
We present the skill-wise breakdown of relative gains of each method discussed in Table 2, in Figure 5A. The largest increases are observed for the language-only skill across all methods, except the image-grounding score which deprioritizes unimodal samples. This result is especially striking because none of the datasets in our experimental setting contain samples similar to those in the MMLU evaluation dataset. It suggests that the underlying LLM in LLaVA can recover this skill from similar multimodal samples. The next highest gains are seen for the knowledge VQA dataset, using our image-grounding score and the Adapt-$\infty$ method. Multilingual skills appear to be the hardest skills to learn and retain over time, potentially because they utilize a different part of the model’s vocabulary than other tasks and are included in the M3IT dataset only.
What is important for multimodal skill recovery?
Multi-way pruning vs. single pruner.
@@
Adapt-$\infty$ uses one among various importance score functions (or pruner experts) for each cluster to select a representative subset of importance samples from the cluster. This works better than using any one single metric across all clusters as shown in Table 3. Importantly, this serves as a flexible framework for swapping or adding advanced ‘expert’ pruners, e.g., learnability score (Mindermann et al., 2022; Evans et al., 2024).
Source of data representations.
@@
Various data selection works propose different ways of representing data samples in a high-dimensional space. Methods designed for pruning pretraining datasets or single-task finetuning can effectively use semantic representations such as sentence embeddings (Abbas et al., 2023) and hidden state outputs (Sorscher et al., 2022; Maharana et al., 2024). However, we observe subpar performance with the use of semantic representations for pseudo-task clustering, as reported in Table 3. Gradient vectors are better representations of the skill component of a data sample as we show in Figure 4. Further, we experimented with gradients from all layers (similar to Xia et al. (2024); Liu et al. (2024)) as well as individual layers of the model. We observed the best performance with gradients from the middle layer only. Gradients from the last layer also work relatively well with Adapt-$\infty$ but those from the first layer work poorly. This discrepancy correlates with the compactness of task clusters in the t-SNE plots of gradients from the corresponding layers, as demonstrated in Figure 5B.
Sampling budget across clusters.
@@
As outlined in Section 4, we sample a subset with an equal number of samples from each pseudo-task cluster in Adapt-$\infty$. The budgets leftover from smaller clusters are distributed equally across the remaining clusters. We experimented with more intuitive budgets for each cluster i.e., based on the density of the cluster members (Abbas et al., 2024). However, it did not result in significant changes in the performance.
Efficiency.
@@
Most data selection methods require extra steps to select an influential data subset and incur computational costs. Adapt-$\infty$ expends additional memory and inference-time compute to effectively select data. We experiment with three methods to reduce this cost: (1) Zero-order gradients (Hinton, 2022), (2) Varying size of data pool in Lite-Adapt-$\infty$, and (3) gradients from a smaller model i.e., TinyLLaVA (Zhou et al., 2024). Using zero-order gradients instead of full gradients leads to a 2% drop in average accuracy and a 2.4% drop in relative gains. TinyLLaVA gradients demonstrate a similar drop, in addition to a higher forgetting rate. Conversely, the performance of Lite-Adapt-$\infty$ improves with increasing size of the data pool post-deduplication (see Table 5).

### v2

# Adapt-$\infty$: Scalable Continual Multimodal Instruction Tuning via Dynamic Data Selection

@@ Abstract
Visual instruction datasets from various distributors are released at different times and often contain a significant number of semantically redundant text-image pairs, depending on their task compositions (i.e., skills) or reference sources. This redundancy greatly limits the efficient deployment of continually adaptable multimodal large language models, hindering their ability to refine existing skills and acquire new competencies over time. To address this, we reframe the problem of lifelong Instruction Tuning (LiIT) via data selection, where the model automatically selects beneficial samples to learn from earlier and new datasets based on the current state of acquired knowledge in the model. Based on empirical analyses that show that selecting the best data subset using a static importance measure is often ineffective for multi-task datasets with evolving distributions, we propose Adapt-$\infty$, a new multi-way and adaptive data selection approach that dynamically balances sample efficiency and effectiveness during LiIT. We first construct pseudo-skill clusters by grouping gradient-based sample vectors. Next, we select the best-performing data selector for each skill cluster from a pool of selector experts, including our newly proposed scoring function, Image Grounding score. This data selector samples a subset of the most important samples from each skill cluster for training. To prevent the continuous increase in the size of the dataset pool during LIT, which would result in excessive computation, we further introduce a cluster-wise permanent data pruning strategy to remove the most semantically redundant samples from each cluster, keeping computational requirements manageable. We validate the effectiveness and efficiency of Adapt-$\infty$ over a sequence of various multimodal instruction tuning datasets with various tasks, including (Knowledge) VQA, multilingual, grounding, reasoning, language-only, and multi-image comprehension tasks. Training with samples selected by Adapt-$\infty$ alleviates catastrophic forgetting, especially for rare tasks, and promotes forward transfer across the continuum using only a fraction of the original datasets. [^1] Code is released at https://github.com/adymaharana/adapt-inf.
@@ 1 Introduction
Multimodal instruction tuning (Liu et al., 2023b; Zhang et al., 2023a; Liu et al., 2023a; Gan et al., 2024; Yoon et al., 2024) has been actively explored to enhance visual reasoning or the generation ability of Multimodal Large Language Models (MLLMs) (Zhu et al., 2023; Li et al., 2023a; Tang et al., 2023; Team et al., 2023; Munasinghe et al., 2023; Yu et al., 2024; Zhang et al., 2024) following user intents by training models on human or machine-generated multi-task visual instruction tuning datasets (Li et al., 2023c; Yin et al., 2023; Xu et al., 2024; Chen et al., 2024). While many distributors continue to release new high-quality instruction-tuning tasks and datasets, continually adapting large models to these massive multi-task datasets over time is prohibitively costly and inefficient. Given a pre-trained MLLM and the continuous expansion of the dataset pool with a stream of multi-task instruction-tuning datasets, as commonly observed in the research community today, the challenge lies in developing an ever-evolving, instruction-following MLLM in the most data- and computation-efficient manner. This research question poses a realistic, sustainable instruction tuning scenario for MLLMs, distinct from conventional continual learning (Zenke et al., 2017; Yoon et al., 2018; Van de Ven & Tolias, 2019), which focuses on learning a sequence of disjoint tasks. Specifically, at each time step, we assume a new multimodal multi-task instruction-tuning dataset is added to the existing training pool, which already contains previous datasets. The model aims to train on samples from this continually expanding dataset pool, a scenario we refer to as Lifelong Instruction Tuning (LiIT). This scenario is critical as recent datasets tend to be exhaustively large to cover a variety of skills in uni- and multimodal tasks. It poses the unique challenges of (1) data redundancies over time due to significant overlap between datasets, (2) imbalanced task representations due to over-representation of simple tasks like captioning, (3) learning rare tasks, and (4) learning new output modalities.
@@
To build an efficient, lifelong-evolving MLLM, we introduce Adapt-$\infty$: Adaptive Multi-way Pruning, an efficient and dynamic multimodal data selection strategy. At each time step, Adapt-$\infty$ selects the most beneficial samples for the current model from the data pool, adapting to the model’s evolving knowledge and changing dataset distributions. This is crucial for LiIT, as sample importance shifts over time. Adapt-$\infty$ operates in two major steps: (1) Task-based clustering. When introducing a new dataset, we integrate it into the training data pool and create pseudo-task clusters using gradient vectors to represent data samples. (2) Cluster-wise data selection. We select influential samples from each cluster for model training. Motivated by the observation that different score functions (Paul et al., 2021; Toneva et al., ; Coleman et al., ) define sample importance differently, we propose a new multi-way data selection approach that chooses the best scoring function from a pool of experts based on its discriminativeness (measured by entropy). It selects a skill-balanced subset of highly influential samples from the current training data pool. In addition, to better assess the influence of multimodal samples, we also propose a new scoring function, image grounding score (IG), which measures the change in sample perplexity when grounded by visual information. This metric prioritizes samples that enhance MLLM multimodal skills and serves as an effective data selector in Adapt-$\infty$. As the data pool grows, computing scores and representations in Adapt-$\infty$, like other data selection strategies, becomes costly. To ensure diversity while managing computational costs, Adapt-$\infty$ adds (3) Permanent data pruning that removes semantically redundant samples from the data pool at the end of each time step, thereby continually controlling its size during LiIT.
@@
We design an experimental setup for this previously unexplored scenario of lifelong multimodal instruction tuning using the pre-trained LLaVA 1.5 (Liu et al., 2023a) model on a stream of five visual instruction tuning datasets. As discussed earlier, experience replay alone is insufficient to mitigate forgetting, however, selecting a random subset from the data pool of past and new datasets reduces the forgetting rate from 26% to nearly 2%. Score-based data-selection strategies largely fail in this setting due to their inability to select task-balanced data subsets from multi-task datasets. Adapt-$\infty$ not only minimizes the forgetting to a mere 0.9% using only a fraction of the training data pool, but also promotes forward transfer of skills and consistently achieves $>$100% relative gains.
@@
We conduct extensive ablations of Adapt-$\infty$ to identify best-performing settings. Our key finding is that hidden layer outputs capture semantics, while gradient vectors represent skills, making them more effective for pseudo-task clustering (see examples in Figures 2 and 4). Further, we find that gradients from the middle layer of the model lead to the best performance, suggesting that skill retention could be localized to a few layers in LLMs. We also find that zero-order gradients (Hinton, 2022; Sung et al., 2024) are promising, computationally cheaper alternatives to backpropagated gradients for pseudo-task clustering (See Table 5). Lastly, skill-wise analysis shows language-only skills are easiest to retain, while multilingual multimodal skills suffer significant forgetting in LiIT.
@@ 2 Related Work
Data selection has been explored in the form of coreset selection in many works (Welling, 2009; Chen et al., 2010; Feldman et al., 2011). Uncertainty/loss/error-based methods estimate the difficulty of a sample from model confidence (Swayamdipta et al., 2020) or its training dynamics (Toneva et al., ; Paul et al., 2021; Bachem et al., 2015). Zheng et al. (2023) address catastrophic accuracy drop at high pruning rates; Maharana et al. (2024) represent datasets as undirected graphs and employ message passing to select the best subset. Gadre et al. (2024) investigate data selection for CLIP Radford et al. (2021) models. Evans et al. (2024) use learnability score (Mindermann et al., 2022) to accelerate training of CLIP models.
Figure 2: A: Sample distributions for different visual language tasks in M3IT (Li et al., 2023b) based on two importance scores, EL2N and entropy. B: t-SNE visualization of sample vectors based on their gradients and features. C: Histogram of Perplexity and Image Grounding scores. We visualize a few samples from M3IT with prompts (black) and ground-truth answers (blue)
@@ 4.4 Combined Permanent-Temporal Data Pruning by Removing Redundancy
The steps of Adapt-$\infty$ outlined in the previous sections enable the model to effectively select the most beneficial samples for each pseudo-task (i.e., skill) and train on them adaptively, based on the current model and evolving dataset distributions. However, as the data pool grows with the release of new instruction-tuning datasets, the computational burden of updating gradient-based sample vectors and ranking their importance increases. To keep the computational requirements manageable over time, we further implement a permanent data pruning strategy to iteratively reduce the size of the entire dataset pool. At the end of training at each timestep, we measure pairwise cosine similarities between samples within each cluster and prune those with maximum similarities, as they are highly redundant and contain overlapping knowledge. The similarity is computed in the semantic space, which is well represented using hidden layer outputs of the model as shown in Figure 4 (Abbas et al., 2023; Sorscher et al., 2022). Larger clusters are prioritized for pruning to fit a predefined data pool budget $D$ until all clusters retain a uniform number of samples. We refer to the version of Adapt-$\infty$ with this combined permanent-temporal pruning as Lite-Adapt-$\infty$.
@@ 5.1 Experimental Setup
We conduct our experiments using the LLaVA 1.5 multimodal large language model (Liu et al., 2023a). It is trained on top of the Vicuna LLM (Chiang et al., 2023) using a corpus of approximately 600K image-text caption pairs for pretraining of the vision projectors from the pre-trained CLIP visual encoder ViT-L/14 (Radford et al., 2021). Further, it is trained on the LLaVA instruction tuning dataset consisting of 665K text-only and vision-language instances. We adopt LoRA finetuning (Hu et al., 2021) of the LLaVA-1.5-7B model with the recommended hyperparameters. [^2] https://github.com/haotian-liu/LLaVA/tree/main
@@
Adapt-$\infty$ Setup. The optimal value of $k$ in the pseudo-task clustering step is computed from a grid search over values of $k$ between 5 and 50, and selected based on the WSS value of clusters. [^3] Within the sum of squares (WSS) is sum of squared distance between samples and their cluster centroids. In the score-based sample selection step, we use a bin size of 50 and discard the top and bottom 5% of samples for computing entropy as well as for CCS sampling, to remove outliers, low-quality samples, and uninformative data (Zheng et al., 2023). We use perplexity, image grounding (Section 3.2), EL2N (Paul et al., 2021) and entropy (Coleman et al., ) score functions for $\mathcal{S}$ throughout the paper.
@@ 5.2 Main Results
Sequential training leads to catastrophic forgetting of skills. Multiple skills learned at $t$=0 (LLaVA) are forgotten at $t$=1,2 upon training the model on datasets containing a different set of skills. The M3IT dataset ($t$=1) does not contain grounding tasks and results in large drops in performance for the same. Similarly, the MiniGPT4 dataset ($t$=2) predominantly contains high-quality captions and causes forgetting of all other skills. The MANTIS dataset ($t$=3) improves performance on the MMMU dataset because it contains training instances with documents, charts, and visualization images. At $t$=4, the LAMM dataset contains object detection and keypoint detection tasks that promote recovery of the referring comprehension skill learned at $t=0$, presumably due to a similarity in their non-textual output modalities. Finally, VisionFLAN ($t$=5) recovers performance on all skills except grounding and MMMU. Surprisingly, VisionFLAN also induces recovery of multilingual and unimodal skills (MMLU) despite not containing those tasks in its dataset composition. This method incurs nearly 26% forgetting across all timesteps and retains only 68% of its all-time high performances at the final timestep (see Table 2). Models trained using Adapt-$\infty$ also outperform other models in multi-turn visual conversations (see Figs. 67 and 8).
@@
The use of gradient vectors in Adapt-$\infty$ to pool samples into pseudo-task clusters before pruning ensures that all skills are well-represented at each time step, resulting in the lowest forgetting rate i.e., 0.9%. Further, the multi-way score-based selection of important samples within those clusters promotes forward transfer of skills and results in $>$100% relative gain, unlike any other method in Table 2. The relative gain increases by nearly 5% on doubling the data budget from 25K to 50K samples and shows signs of plateauing with further increase in data.
@@ 6 Analysis
Relative gains per skill. We present the skill-wise breakdown of relative gains of each method discussed in Table 2, in Figure 5A. The largest increases are observed for the language-only skill across all methods, except the image-grounding score which deprioritizes unimodal samples. This result is especially striking because none of the datasets in our experimental setting contain samples similar to those in the MMLU evaluation dataset. It suggests that the underlying LLM in LLaVA can recover this skill from similar multimodal samples. The next highest gains are seen for the knowledge VQA dataset, using our image-grounding score and the Adapt-$\infty$ method. Multilingual skills appear to be the hardest skills to learn and retain over time, potentially because they utilize a different part of the model’s vocabulary than other tasks and are included in the M3IT dataset only.
@@
Multi-way pruning vs. single pruner. Adapt-$\infty$ uses one among various importance score functions (or pruner experts) for each cluster to select a representative subset of importance samples from the cluster. This works better than using any one single metric across all clusters as shown in Table 3. Importantly, this serves as a flexible framework for swapping or adding advanced ‘expert’ pruners, e.g., learnability score (Mindermann et al., 2022; Evans et al., 2024).
@@
Source of data representations. Various data selection works propose different ways of representing data samples in a high-dimensional space. Methods designed for pruning pretraining datasets or single-task finetuning can effectively use semantic representations such as sentence embeddings (Abbas et al., 2023) and hidden state outputs (Sorscher et al., 2022; Maharana et al., 2024). However, we observe subpar performance with the use of semantic representations for pseudo-task clustering, as reported in Table 3. Gradient vectors are better representations of the skill component of a data sample as we show in Figure 4. Further, we experimented with gradients from all layers (similar to Xia et al. (2024); Liu et al. (2024)) as well as individual layers of the model. We observed the best performance with gradients from the middle layer only. Gradients from the last layer also work relatively well with Adapt-$\infty$ but those from the first layer work poorly. This discrepancy correlates with the compactness of task clusters in the t-SNE plots of gradients from the corresponding layers, as demonstrated in Figure 5B.
@@
Sampling budget across clusters. As outlined in Section 4, we sample a subset with an equal number of samples from each pseudo-task cluster in Adapt-$\infty$. The budgets leftover from smaller clusters are distributed equally across the remaining clusters. We experimented with more intuitive budgets for each cluster i.e., based on the density of the cluster members (Abbas et al., 2024). However, it did not result in significant changes in the performance.
@@
Efficiency. Most data selection methods require extra steps to select an influential data subset and incur computational costs. Adapt-$\infty$ expends additional memory and inference-time compute to effectively select data. We experiment with three methods to reduce this cost: (1) Zero-order gradients (Hinton, 2022), (2) Varying size of data pool in Lite-Adapt-$\infty$, and (3) gradients from a smaller model i.e., TinyLLaVA (Zhou et al., 2024). Using zero-order gradients instead of full gradients leads to a 2% drop in average accuracy and a 2.4% drop in relative gains. TinyLLaVA gradients demonstrate a similar drop, in addition to a higher forgetting rate. Conversely, the performance of Lite-Adapt-$\infty$ improves with increasing size of the data pool post-deduplication (see Table 5).
See Appendix C for additional results using LLMs, analysis on efficiency and recovery of skills.